BRICS Notes Series, 2004

September 16, 2004

This document is also available as PostScript and DVI.

References

[NS-04-1] Abstract, PDF.

Luca Aceto, Willem Jan Fokkink, and Irek Ulidowski, editors. Preliminary Proceedings of the Workshop on Structural Operational Semantics, SOS ’04, (London, United Kingdom, August 30, 2004), August 2004. vi+56.